OpenStudy (anonymous):

Solve the absolute value equations below:

OpenStudy (anonymous):

\[|x + 6|= 3\] A) x = 3 B) x = 3, -3 C) x = -3, -9 D) x = 6, -6

OpenStudy (igreen):

This breaks into two equations: \(x + 6 = 3\) and \(x - 6 = 3\) ----------------- \(x+ 6 = 3\) Subtract 6 to both sides: \(x = -3\) ----------------- \(x - 6 = 3\) Add 6 to both sides: What's 6 + 3? @pigglet

OpenStudy (anonymous):

9

OpenStudy (igreen):

Yep, so the solutions are -3, and 9.

OpenStudy (anonymous):

it would be -3,-9. there is no -3,9

OpenStudy (anonymous):

The second equation is -x-6 not x-6 I believe.

OpenStudy (anonymous):

As it is the negative form of (x+6)

OpenStudy (igreen):

Oh..that's right, my bad..I made the second equation wrong. So it should be (-3, -9). 'Cause you have to multiply -1 to both sides which gives you -9..

OpenStudy (igreen):

Thanks @alrightatmaths

OpenStudy (anonymous):

So its C

OpenStudy (anonymous):

That's ok, no worries

OpenStudy (igreen):

Yes. @pigglet